Database last updated on Wed, 28 Jan 2026 09:40:47 +0000 by localhost

Instrument constant changes for the week before 2025012 (3 sec, 0 rows)

HG burned out? (AP results when HG is ON) (0 sec, 1 rows)

Measurements and tests summary.


_jday_loc_serial_skc
012LONDON#019menu
(0 sec)

Dead time (0 sec)

Nothing to report.

Standard lamp R6 (1 sec)

Nothing to report.

Standard lamp R5 (1 sec)

Nothing to report.

Number of days since last dome/window cleaning (0 sec, 2 rows)

serialdays
019more than a year

Missing files (0 sec, 1 rows)

Seriallast datadays ago
01920250130-1

Run/Stop results for last week. (0 sec, 1 rows)

SerialTotalFailed 0Failed darkFailed 1Failed 2Failed 3Failed 4Failed 5Failed double
0192601000011

Ozone results for good DS obs (std<2.5DU, mu<4) within 3 minutes between 2 Brewers. (1 sec)

No ozone measurements to display.

Ozone results for good FM obs (std<5DU, mu<4) within 3 minutes between 2 Brewers. (0 sec)

No ozone measurements to display.

Unsorted comments (0 sec, 9424 rows)

SerialtypeTextCountYYYYJJJ
019env:DosBox12025012
019cubdsp:012025012

The END (0 sec, 2 rows)

If you have questions and/or comments about this page please contact Volodya Savastiouk